### v1

# Uncertainty Quantification and Confidence Calibration in Large Language Models: A Survey

Abstract.
Large Language Models (LLMs) excel in text generation, reasoning, and decision-making, enabling their adoption in high-stakes domains such as healthcare, law, and transportation. However, their reliability is a major concern, as they often produce plausible but incorrect responses. Uncertainty quantification (UQ) enhances trustworthiness by estimating confidence in outputs, enabling risk mitigation and selective prediction. However, traditional UQ methods struggle with LLMs due to computational constraints and decoding inconsistencies. Moreover, LLMs introduce unique uncertainty sources, such as input ambiguity, reasoning path divergence, and decoding stochasticity, that extend beyond classical aleatoric and epistemic uncertainty. To address this, we introduce a new taxonomy that categorizes UQ methods based on computational efficiency and uncertainty dimensions (input, reasoning, parameter, and prediction uncertainty). We evaluate existing techniques, assess their real-world applicability, and identify open challenges, emphasizing the need for scalable, interpretable, and robust UQ approaches to enhance LLM reliability.
Uncertainty Quantification; Large Language Models
copyright: none
1. Introduction
Large Language Models (LLMs) like GPT-4 (Achiam et al., 2023) and PaLM (Anil et al., 2023) have achieved remarkable capabilities in text generation, reasoning, and decision-making, driving their adoption in high-stakes domains such as healthcare diagnostics (Qiu et al., 2024; Ullah et al., 2024; Goh et al., 2024; Da et al., 2024d), legal analysis (Cheong et al., 2024; Jayakumar et al., 2023; Drápal et al., 2023), and transportation systems (Jin et al., 2023; Da et al., 2024b; Zheng et al., 2023a; Lai et al., 2023). However, their reliability remains a critical concern: LLMs often produce plausible but incorrect or inconsistent outputs, with studies showing that over 30% of answers in medical QA tasks contain factual errors (Jin et al., 2021). In sensitive applications, these limitations pose risks ranging from misinformation to life-threatening misdiagnoses, underscoring the urgent need for robust reliability frameworks.
Uncertainty quantification (UQ) emerges as a pivotal mechanism to enhance LLM trustworthiness by explicitly modeling confidence in model outputs. By estimating uncertainty, users can identify low-confidence predictions for human verification, prioritize high-certainty responses, and mitigate risks like overconfidence in hallucinations (Slobodkin et al., 2023; Liu et al., 2024a). For instance, in clinical settings, uncertainty-aware LLMs could flag uncertain diagnoses for specialist review, reducing diagnostic errors by up to 41% (Sen et al., 2024). This capability is particularly critical as LLMs transition from experimental tools to production systems requiring accountability.
Traditional UQ methods face significant hurdles when applied to Large Language Models (LLMs). Bayesian approaches like Monte Carlo dropout (Gal and Ghahramani, 2016) are computationally prohibitive for trillion-parameter models and natural language generation (NLG) tasks, while ensemble methods struggle with consistency across diverse decoding strategies (Lu et al., 2024). Furthermore, LLMs introduce unique uncertainty sources—such as input ambiguity, reasoning path divergence, and decoding stochasticity—that transcend classical aleatoric and epistemic categorizations (Hüllermeier and Waegeman, 2021). The complexity of LLMs, characterized by sequence generation over vast parameter spaces and reliance on massive datasets, exacerbates uncertainty challenges. This complexity, coupled with the critical need for reliable outputs in high-stakes applications, positions UQ for LLMs as a compelling yet underexplored research frontier.
This tutorial introduces a novel taxonomy for LLM uncertainty quantification, categorizing methods along two axes: (1) computational efficiency (e.g., single-pass vs. sampling-based techniques) and (2) uncertainty dimensions (input, reasoning, parametric, predictive). Our framework addresses three gaps in prior work: First, it decouples uncertainty sources unique to LLMs from traditional ML contexts. Second, it evaluates methods through the lens of different dimensions of the responses from LLM: input uncertainty, reasoning uncertainty, parameter uncertainty, and prediction uncertainty. Each of these dimensions may involve aleatoric uncertainty, epistemic uncertainty, or a mixture of both. Third, it identifies understudied areas like reasoning uncertainty, which accounts for 58% of errors in multi-step QA tasks (Da et al., 2025a).
Contributions: We provide (1) the first systematic review of UQ methods tailored to different dimensions of uncertainty in LLM, including input, reasoning, parameter, and prediction uncertainty; (2) a comprehensive introduction to the evaluation of UQ methods and the challenges.
Connection to Existing Surveys: Prior surveys (Shorinwa et al., 2024; Huang et al., 2024b; Beigi et al., 2024; Huang et al., 2025a; Rawte et al., 2023; Zhang et al., 2023) focus narrowly on hallucination detection or retrofitting classical UQ taxonomies, neglecting LLM-specific challenges like prompt-driven input uncertainty. Our work uniquely addresses the interplay between model scale, open-ended generation, and uncertainty dynamics—factors critical for modern LLMs but overlooked in earlier frameworks.
The remainder of this survey is structured as follows: Section 2 characterizes LLM uncertainty dimensions and differentiates confidence from uncertainty. Section 3 evaluates UQ methods using our taxonomy. Section 4 introduces the evaluation of UQ methods for LLM. Sections 5 and 6 introduce the applications of UQ in different domains with LLMs and identify open challenges and future directions. Table 1 shows the notation and Table 2 shows an overview of all methods we discuss in this paper.
2. Perliminaries
2.1. Sources of Uncertainty in LLMs
2.1.1. Aleatoric vs. Epistemic Uncertainty
For uncertainty quantification of traditional machine learning tasks such as classification or regression (Young et al., 2024), there are mainly two types of uncertainty (Der Kiureghian and Ditlevsen, 2009): aleatoric and epistemic uncertainty. Aleatoric uncertainty defines the uncertainty from noise in the dataset. Epistemic uncertainty, on the other hand, arises from the model’s lack of knowledge about the underlying data distribution.
Aleatoric uncertainty in LLMs primarily stems from data sources used to train LLMs, which contain inconsistencies, biases, and contradicting information. Furthermore, ambiguity in natural language contributes to aleatoric uncertainty, as different interpretations of the same prompt can lead to multiple plausible responses.
On the other hand, When encountering unfamiliar topics, LLMs may exhibit high epistemic uncertainty, often manifesting as hallucinations or overconfident yet incorrect statements. Epistemic uncertainty can be reduced through domain-specific fine-tuning or retrieval-augmented generation techniques that allow the model to access external knowledge sources.
Though uncertainty for LLMs can also be classified as these two categories, these two categories alone are insufficient to fully capture the complexities of uncertainty in the language model. The unique nature of LLM inference introduces additional uncertainty factors. In particular, LLMs exhibit uncertainty not only due to training data limitations but also due to input variability and decoding mechanisms. Therefore, to address these challenges, we introduce four new dimensions of uncertainty.
2.1.2. Uncertainty with Different Dimensions
To systematically analyze uncertainty in LLMs, we categorize it into four key dimensions: input uncertainty, reasoning uncertainty, parameter uncertainty, and prediction uncertainty. Each dimension may involve aleatoric uncertainty, epistemic uncertainty, or a combination of both. This structured framework provides a more comprehensive understanding of uncertainty quantification in LLMs.
Input Uncertainty (Aleatoric Uncertainty): Input uncertainty arises when a prompt is ambiguous or underspecified, making it impossible for an LLM to generate a single definitive response. This is inherently aleatoric, as even a “perfect model” cannot resolve the ambiguity. For instance, “What is the capital of this country?” lacks sufficient context, leading to unpredictable outputs. Similarly, “Summarize this document” may yield different responses depending on different expected details.
Reasoning Uncertainty (Mixed Uncertainty): Reasoning uncertainty occurs when an LLM derives answers through multi-step logic or retrieval, leading to ambiguous or incorrect reasoning. This uncertainty is aleatoric when the problem itself is ambiguous and epistemic when the model cannot offer robust reasoning.
Parameter Uncertainty (Epistemic Uncertainty): Parameter uncertainty stems from gaps in the training data, where the model has either never seen relevant information or has learned an incorrect representation. Unlike aleatoric uncertainty, epistemic uncertainty can be reduced by improving the model’s knowledge base. Bayesian methods (Gal and Ghahramani, 2016), deep ensembles (Lakshminarayanan et al., 2017), and uncertainty-aware training (Mukherjee and Awadallah, 2020) can help quantify and mitigate this type of uncertainty.
Prediction Uncertainty(Mixed Uncertainty): Prediction uncertainty refers to variability in generated outputs across different sampling runs, influenced by both aleatoric and epistemic sources. For example, when asked, “What are the side effects of a new experimental drug?” the model’s responses might vary significantly across different sampling runs, especially if no reliable data is available in its training set. A high-variance output distribution in such scenarios suggests that the model is both aware of multiple possible answers, reflecting aleatoric uncertainty, and uncertain due to incomplete knowledge, highlighting epistemic uncertainty.
Notation Description
$x$ The question that LLMs answer
$s$ Generation from LLMs
$w_{i}$ i-token in the generation $s$
$\mathcal{D}$ Dictionary of LLMs
$U(x)$ Uncertainty of question $x$
$C(x,s)$ Confidence of generation $s$ given $x$
$H(s)$ Entropy of generation $s$
Table 1. Notation used in this paper.
2.2. Uncertainty and Confidence in LLMs
Uncertainty quantification and confidence estimation are closely related yet distinct concepts in the context of large language models (LLMs). Uncertainty is a property of the model’s predictive distribution, capturing the degree of variability or unpredictability given a particular input. In contrast, confidence reflects the model’s belief in the correctness of a particular answer or prediction. As a concrete example, in the context of classification, a simple confidence measure is the predicted probability $\hat{p}(Y=y|x)$ (an uncertainty measure which does not depend on the particular prediction $y$ could be entropy, taking the form of $\sum_{y}-\hat{p}(Y=y|x)\log{\hat{p}(Y=y|x)}$). The corresponding confidence score in NLG (for an auto-regressive LM) is the joint probability for the generated sequence:
(1) $C(\mathbf{x},\mathbf{s})=\hat{p}(\mathbf{s}|\mathbf{x})$ $=\prod_{i}\hat{p}(s_{i}|\mathbf{s}_{<i},\mathbf{x}).$
The log of Eq. 1 is sometimes referred to as sequence likelihood (Zhao et al., 2023a). In general, while an uncertainty estimate takes the form of $U(x)$, confidence estimates could be expressed as $C(\mathbf{x},\mathbf{s})$. Note that unlike classification tasks, not all NLG applications have the notion of a “correct” answer (e.g. summarization). Thus, while for the ease of writing we use the term correctness throughout this section, it should really be interpreted as the gold-label for the particular application. Note also that in most cases, the correct answer is not unique, and thus such gold-label typically takes the form of a “correctness function” that decides whether a particular generation $\mathbf{s}$ is good or not. We will denote such a function as $f(\mathbf{s}|\mathbf{x})$.
There are usually two dimensions along which researchers improve confidence estimates in NLG, which is unsurprisingly largely influenced by confidence scoring literature from classification (Jiang et al., 2018; Corbière et al., 2019), especially binary classification. We refer to them as ranking performance and calibration:
  • Ranking performance refers to the discriminative power of the confidence measure on the correctness. Like in classification, LLM confidence is often evaluated by its ability to separate correct and incorrect answers, thus typically measured by evaluation metrics like AUROC (Kadavath et al., 2022) or AUARC (Lin et al., 2024) as detailed in Section 4.
  • Calibration refers to closing the gap between the confidence score and the expected correctness conditioned on confidence score. It has a long history preceding LLM or even modern machine learning (Murphy and Winkler, 1977; Degroot and Fienberg, 1983; Murphy and Winkler, 1984), but bears slightly different meanings in NLP. In general, we could define a perfectly calibrated confidence measure to achieve the following:
    (2) $\forall c,\mathbb{E}[f(\mathbf{s}|\mathbf{x})|C(\mathbf{x},\mathbf{s})=c]=c,$
    where the expectation is taken over the joint distribution of $\mathbf{x}$ and generation $\mathbf{s}$. A lot of papers focus on evaluating the calibration quality of specific LMs and tasks (Wang et al., 2020; Khojah et al., 2022; Kumar and Sarawagi, 2019). Evaluation typically relies on variants of Expected Calibration Error (ECE) (Tian et al., 2023; Kumar and Sarawagi, 2019; Li et al., 2024b). Oftentimes confidence scores from classification could be directly applied (Stengel-Eskin and Van Durme, 2023; Zhao et al., 2021; Jiang et al., 2021) in order to evaluate whether an LM is over- or under-confident, especially for de facto classification tasks like sentiment analysis or multiple-choice QA.
As uncertainty and confidence are often intertwined, many approaches used in uncertainty quantification have their counterpart in confidence estimation. For example, for black-box methods, Lin et al. (2023) computes a similarity matrix of sampled responses and derives confidence estimates for each generation via its degree or distance derived from the graph Laplacian, before using these scores to compute uncertainty. Zhang et al. (2024a) extends such black-box methods to longer generations. For logit-based methods, Malinin and Gales (2021) normalize Eq. 1 with the length of $\mathbf{s}$. Further improvements include replacing the logit-sum or mean with weighted sum, by attention values on the important tokens (Lin et al., 2024) or by importance inferred from natural language inference (NLI) models (Duan et al., 2024). Such variants of sequence likelihood could then be fed for (entropy-style) uncertainty computation (Kuhn et al., 2023; Lin et al., 2024).
Another popular approach is asking the LM itself whether a particular free-form generation is correct (Kadavath et al., 2022). However, this formulation also poses a restriction on the confidence estimation method, as it is essentially a scalar logit. Thus, many extensions focus on applying calibration methods from classification to calibrate such self-evaluation. The few exceptions include Ren et al. (2023b), which generalizes Kadavath et al. (2022) and converts samples from free-form generation into a multiple-choice question (with generations being the options) and adding a ”None of the above” option to elicit the confidence. Similarly, Shrivastava et al. (2025) performs the explicit relative comparison of pairs of answers and then aggregates the preferences into an absolute confidence level.
Since we typically care about the LM’s confidence in the “semantic space” due to semantic invariance, instead of manipulating logits, a popular approach is to perform additional training for confidence estimation. This could be done on the base LM (either full LM (Zhao et al., 2023a; Jiang et al., 2021; Kapoor et al., 2024) or partial (Liu et al., 2024a)) with a different loss, or using a separate model on the internal or external representations from the base LM (Jagannatha and Yu, 2020; Mielke et al., 2022; Ren et al., 2023a; Azaria and Mitchell, 2023; Ulmer et al., 2024). On the other end of the spectrum, without any training, prompting could be used to elicit verbalized confidence values (Xiong et al., 2024a; Tian et al., 2023), or to recalibrate LLM confidence for a particular distribution (Li et al., 2024b) via in-context learning. Finally, one could combine multiple confidence estimation methods and enjoy the benefit of ensembling (Gao et al., 2024).
Just like the evaluation for uncertainty quantification (more in Section 4), the choice of correctness function has a profound impact on the conclusion of the experiments, especially for free-form generation tasks. Popular choices include using (potentially larger) LLM as judges (Liu et al., 2024a; Lin et al., 2023; Tian et al., 2023), human annotations (Ren et al., 2023b; Mielke et al., 2022), or lexical similarities such as ROUGE (Kuhn et al., 2023; Zhao et al., 2023a). Recently, Liu et al. (2025) proposes to evaluate free-form generation confidence measures with selected multiple-choice datasets as an efficient complement. For longer generations, Huang et al. (2024a) proposes to use ordinal (not binary) correctness values to capture the ambiguity in the quality of a generation. In a similar flavor, Baan et al. (2022) studies the issues in the evaluation of calibration (against the human majority) when there is intrinsic human disagreement on the label.
Remarks. Existing literature sometimes uses the terms uncertainty and confidence interchangeably. They do often seemingly coincide: When a model’s prediction has low confidence, we naturally consider this as a high uncertainty case. This, however, is treating $U(\mathbf{x})=-\max_{\mathbf{s}}C(\mathbf{x},\mathbf{s})$ as an uncertainty estimate. In general, a model may exhibit high uncertainty over its output space but still express high confidence in a specific output. Conversely, a model could have low overall uncertainty but low confidence in a particular prediction. While the “low uncertainty low confidence case” is relatively less interesting in classification or regression tasks due to MLE point prediction, this scenario is notably more common in NLG, as the output is typically randomly sampled [^1] In fact, even if the output is greedily generated, it might not have the highest confidence as measured by Eq. 1. from the predictive distribution. There are also applications that require one but not the other (e.g. conformal language modeling (Quach et al., 2023) or seletive generation (Cole et al., 2023)). In the rest of this paper, we sometimes follow the language of the original papers and treat confidence estimates as uncertainty, but will clearly mark the methods that provide confidence estimates.
3. UQ Methods for Different Dimensions
Method Uncertainty Stages Efficency Features Access to model Confidence
Input clarification ensembles (Hou et al., 2024) Input Uncertainty Multi Rounds Generations Black-box No
Uncertainty Quantification for In-Context Learning (Ling et al., 2024) Input Uncertainty Multi Rounds Generations Black-box No
SPUQ (Gao et al., 2024) Input Uncertainty Multi Rounds Generations + Additional Model Black-box No
UAG (Yin et al., 2024) Reasoning Uncertainty Single Round Generation White-box No
CoT-UQ (Zhang and Zhang, 2025) Reasoning Uncertainty Single Round Generation White-box Yes
TouT (Mo and Xin, 2024) Reasoning Uncertainty Multi Rounds Generations Black-box No
TopologyUQ (Da et al., 2025a) Reasoning Uncertainty Multi Rounds Generations Black-box No
Stable Explanations Confidence (Becker and Soatto, 2024) Reasoning Uncertainty Multi Rounds Generation Black-box Yes
SAPLMA (Azaria and Mitchell, 2023) Parameter + Prediction Uncertainty Fine-tuning White-box Yes
Supervised uncertainty estimation(Liu et al., 2024c) Parameter + Prediction Uncertainty Fine-tuning White-box Yes
UaIT (Liu et al., 2024b) Parameter + Prediction Uncertainty Fine-tuning White-box Yes
LoRA ensembles (Balabanov and Linander, 2024) Parameter Uncertainty Fine-tuning White-box Yes
BloB (Wang et al., 2025) Parameter Uncertainty Fine-tuning White-box Yes
BLoRA (Yang et al., 2023) Parameter Uncertainty Fine-tuning White-box Yes
Perplexity (Mora-Cross and Calderon-Ramirez, 2024; Margatina et al., 2023) Prediction Uncertainty Single Round Generation White-box Yes
Shifting Attention to Relevance (SAR) (Duan et al., 2024) Prediction Uncertainty Single Round Generation White-box Yes
P(True) (Kadavath et al., 2022) Prediction Uncertainty Single Round Generation White-box Yes
Response improbability (Fadeeva et al., 2024) Prediction Uncertainty Single Round Generation White-box Yes
Average log probability (Manakul et al., [n. d.]) Prediction Uncertainty Single Round Generation White-box Yes
Predictive Entropy (Kadavath et al., 2022) Prediction Uncertainty Multi Rounds Generations White-box Yes
Relative Mahalanobis distance (Ren et al., 2023a) Prediction Uncertainty Multi Rounds Generations White-box Yes
HUQ (Vazhentsev et al., 2023) Prediction Uncertainty Multi Rounds Generations White-box Yes
Conformal Prediction (Quach et al., 2023; Kumar et al., 2023) Prediction Uncertainty Multi Rounds Generations White-box No
ConU (Wang et al., 2024a) Prediction Uncertainty Multi Rounds Generations White-box No
Level-adaptive conformal prediction (Cherian et al., 2025) Prediction Uncertainty Multi Rounds Generations White-box No
LoFreeCP (Su et al., 2024) Prediction Uncertainty Multi Rounds Generations Black-box No
Ecc(J),Deg(J) (Lin et al., 2023) Prediction Uncertainty Multi Rounds Generations Black-box Yes
Eig(J) (Lin et al., 2023) Prediction Uncertainty Multi Rounds Generations Black-box No
Normal length predictive entropy (Malinin and Gales, 2021) Prediction Uncertainty Multi Rounds Generations +Additional Model White-box Yes
Semantic Entropy (Kuhn et al., 2023) Prediction Uncertainty Multi Rounds Generations + Additional Model White-box Yes
Kernel Semantic Entropy (Nikitin et al., 2024) Prediction Uncertainty Multi Rounds Generations + Additional Model White-box Yes
Ecc(C),Ecc(E),Deg(C),Deg(E) (Lin et al., 2023) Prediction Uncertainty Multi Rounds Generations + Additional Model Black-box Yes
Eig(C),Eig(E) (Lin et al., 2023) Prediction Uncertainty Multi Rounds Generations + Additional Model Black-box No
MD-UQ (Chen et al., 2025) Prediction Uncertainty Multi Rounds Generations + Additional Model Black-box No
U-UE (Da et al., 2024a) Prediction Uncertainty Multi Rounds Generations + Additional Model Black-box Yes
Table 2. An overview of uncertainty quantification methods discussed in this paper.
3.1. Input Uncertainty
As mentioned in Section 2.1.2, input uncertainty arises from the ambiguous or incomplete input to the LLMs and many works in the LLMs domain try to deal with ambiguity. For example, there are many datasets on ambiguity with different tasks such as ambiguity detection (Min et al., 2020a; Guo et al., 2021) and many works on dealing with ambiguity (Zamani et al., 2020; Deng et al., 2023). However, these works did not consider uncertainty at all.
As far as we know, most uncertainty quantification methods that care about input uncertainty focus on perturbing the input prompts of LLMs. For instance, Hou et al. (2024) proposes an approach that generates multiple clarifications for a given prompt and ensembles the resulting generations by using mutual information to capture the disagreement among the predictions arising from different clarifications. Similarly, Ling et al. (2024) quantified the input uncertainty in the setting of in-context learning by using different in-context samples. Gao et al. (2024) proposes SPUQ, which perturbs the input by techniques such as paraphrasing and dummy tokens to expose the model’s sensitivity and capture uncertainty. Specifically, SPUQ quantified the input uncertainty by using a similarity metric such as BERTScore (Zhang et al., 2019) to measure how consistent the responses are across different perturbations.
We can see that there are only a few papers that care about input uncertainty and its application to ambiguity. Since ambiguity is common and important in natural language, we call more effort into input uncertainty and its application.
3.2. Reasoning Uncertainty
Recent research has advanced stepwise uncertainty quantification in LLM reasoning by explicitly eliciting and analyzing the internal reasoning process. For example, Tree of Uncertain Thoughts (TouT) (Mo and Xin, 2024) extends the Tree of Thoughts (ToT) (Yao et al., 2023) framework by addressing the inherent local uncertainties that occur during intermediate reasoning steps. TouT leverages Monte Carlo Dropout to assign uncertainty scores to pivotal decision points and, by integrating these local measures with global search techniques, it enhances the precision of response generation. Similarly, TopologyUQ (Da et al., 2025a) introduces a formal method to extract and structure LLM explanations into graph representations, quantifying reasoning uncertainty by employing graph-edit distances and revealing redundancy through stable topology measures. In addition, Uncertainty-aware Adaptive Guidance (UAG) (Yin et al., 2024) tackles error accumulation in multi-step reasoning by monitoring the predicted probability of the next token at each generation step, dynamically retracting to more reliable states and incorporating certified reasoning clues when high uncertainty is detected. Complementing these methods, Stable-Explanation Confidence (Becker and Soatto, 2024) quantifies uncertainty by examining the distribution of generated explanations, treating each model’s explanation pair as a test-time classifier to construct a posterior answer distribution that reflects overall reasoning confidence. More recently, CoT-UQ (Zhang and Zhang, 2025) has integrated chain-of-thought reasoning into a response-level uncertainty quantification framework, thereby leveraging the inherent multi-step reasoning capability of LLMs to further improve uncertainty assessment. Collectively, these approaches provide a robust and interpretable framework for enhancing LLM reasoning by quantifying uncertainty at both local and global levels.
3.3. Parameter Uncertainty
Parameter uncertainty arises when an LLM lacks sufficient knowledge due to limitations in its training data or model parameters. Parameter uncertainty reflects the model’s uncertainty about its own predictions, which can be reduced with additional training or better adaptation techniques.
Traditional UQ methods like Monte Carlo (MC) Dropout and Deep Ensembles have been widely used but are computationally infeasible for large-scale LLMs due to the need for multiple forward passes or model replicas. To address this, Bayesian Low-Rank Adaptation by Backpropagation (BLoB)(Wang et al., 2025) and Bayesian Low-Rank Adaptation (BLoRA)(Yang et al., 2023) incorporate Bayesian modeling into LoRA adapters, allowing uncertainty estimation through parameter distributions without full-model ensemble. However, these methods still incur significant computational costs.
Finetuning-based approaches offer a more practical alternative. Techniques such as Supervised Uncertainty Estimation(Liu et al., 2024c) train auxiliary models to predict the confidence of LLM outputs based on activation patterns and logit distributions. Similarly, Uncertainty-aware Instruction Tuning (UaIT)(Liu et al., 2024b) modifies the fine-tuning process to explicitly train models to express uncertainty in their outputs. SAPLMA(Azaria and Mitchell, 2023) refines probabilistic alignment techniques to dynamically adjust model uncertainty estimates, ensuring adaptability to different downstream tasks. Additionally, LoRA ensembles(Balabanov and Linander, 2024) provide an alternative to full-model ensembles by training multiple lightweight LoRA-adapted variants of an LLM instead of retraining the entire network.
3.4. Prediction Uncertainty
Most off-the-shelf uncertainty quantification methods focus on prediction uncertainty since it is the most straightforward way to estimate the uncertainty.
3.4.1. Single Round Generation
Most single-round generation methods utilize the logit or hidden states during the generation procedure. Single-round generation methods have the highest efficiency.
Perplexity is a measure of how well a probabilistic language model predicts a sequence of text (Vaswani et al., 2017) while Mora-Cross and Calderon-Ramirez (2024), Margatina et al. (2023) and Manakul et al. ([n. d.]) utilize the perplexity as the uncertainty. In detail, using $w_{i}$ as the i-th token in the generation, perplexity is given by:
(3) $\text{Perplexity}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\ln p(w_{i})\right)$
A higher perplexity means the model spreads its probability more broadly over possible words, indicating that it has a higher uncertainty.
Maximum Token Log-Probability. Apart from the perplexity, Manakul et al. ([n. d.]) also uses maximum token log-probability:
(4) $Max(p)=\max\limits_{i}(-\ln p(w_{i}))$
Maximum token log-probability measures the sentence’s likelihood by assessing the least likely token in the sentence. A higher $\text{Maximum}(p)$ indicates higher uncertainty of the whole generation.
Entropy reflects how widely distributed a model’s predictions are for a given input, indicating the level of uncertainty in its outputs (Malinin and Gales, 2018; Kuhn et al., 2023; Kadavath et al., 2022).Entropy for the i-th token is provided by:
(5) $\mathcal{H}_{i}=-\sum_{\tilde{w}\in\mathcal{D}}p_{i}(\tilde{w})\log p_{i}(\tilde{w})$
Then it is possible to use the mean or maximum value of entropy as the final uncertainty (Manakul et al., [n. d.]):
(6) $Avg(\mathcal{H})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{H}_{i};Max(\mathcal{H})=\max\limits_{i}(\mathcal{H}_{i})$
Furthermore, Shifting Attention to Relevance (SAR), proposed by Duan et al. (2024), enhanced the performance of entropy by adjusting attention to more relevant tokens inside the sentence. In detail, SAR assigned weight for $\mathcal{H}_{i}$ and the weight $R(w_{i},s,x)$ can be obtained by:
(7) $R(w_{i},s,x)=1-\left|g(x\cup s,x\cup s\setminus\{w_{i}\})\right|,$
Where $g$ is a function that measures the semantic similarity between two sentences such as NLI models.
Response Improbability. Fadeeva et al. (2024) uses response improbability, which computes the probability of a given sentence and subtracts the resulting value from one. In detail, response improbability is provided by:
(8) $MP(s)=1-\prod_{i=1}p_{i}(w_{i}).$
If the sentence is very certain (i.e., the product of token probabilities is high), $MP(s)$ will be low.
P(True) Kadavath et al. (2022) proposes P(True) [^2] The original name is P(IK), which stands for “I Know”., which measures the uncertainty of the claim by asking the LLM itself whether the generation is true or not. Specifically, P(True) is calculated:
[…]
Note that here we are using $y_{1}$ as the first token instead of $w_{1}$ because $w_{1}$ represents the first token to the generation $s$ while $y_{1}$ represented the first token when asking LLM whether the generation $s$ is correct or not. P(True) requires to run the LLM twice. However, it does not require multiple generations $s$. Therefore, we still classify this method to single-round generation. (This could be considered an uncertainty estimate as the sequence to be evaluated is the prediction given the input.)
3.4.2. Multiple rounds generation
Multiple rounds generation methods estimate uncertainty by generating multiple predictions from the LLMs and analyzing their consistency, similarity, or variability. These approaches assume that if a model is confident, its outputs should be stable across different sampling conditions.
Token-Level Entropy. Token-level entropy quantifies uncertainty in LLMs by analyzing the probability distribution of generated tokens across multiple samples. A confident model assigns high probability to a specific token, resulting in low entropy, while uncertain predictions distribute probability across multiple tokens, leading to higher entropy.
Multiple responses are generated for the same input to estimate token-level entropy, and the entropy of the token probability distribution is computed. For example, predictive entropy (Kadavath et al., 2022) can also be applied to multiple response settings and shows a better uncertainty quality based on multiple outputs variability. Similarly, SAR (Duan et al., 2024) could also be applied to multiple responses. Another study refines this method with Monte Carlo-based approximations (Malinin and Gales, 2021). It focuses on how probability distributions evolve across tokens during autoregressive generation. There are two main approaches: one that directly estimates uncertainty by averaging entropy across multiple sampled outputs and another that decomposes sequence-level uncertainty into token-level contributions using a structured entropy approximation.
Conformal Prediction. Conformal Prediction (CP) is a statistical framework that provides formal coverage guarantees for uncertainty estimation in LLMs. It has been widely adopted in recent research due to its distribution-free properties, making it suitable for both black-box and white-box models.
In the black-box setting, where model internals are inaccessible, CP estimates uncertainty using response frequency, semantic similarity, or self-consistency. One study proposes a method tailored for API-only LLMs (Su et al., 2024), using frequency-based sampling combined with normalized entropy (NE) and semantic similarity (SS) to define nonconformity scores. Another black-box CP method introduces a self-consistency-based uncertainty measure (Wang et al., 2024a), which clusters sampled generations and selects a representative response to construct prediction sets with correctness guarantees, making it particularly effective for open-ended NLG tasks.
On the other hand, white-box CP methods use logits, internal activations, and calibration techniques for more refined uncertainty estimation. One study proposes Conformal Language Modeling (Quach et al., 2023), which integrates CP with autoregressive text generation by dynamically calibrating a stopping rule to ensure at least one response in the generated set is statistically valid. Another work adapts CP for multiple-choice QA (Kumar et al., 2023), using model confidence scores to calibrate prediction sets, ensuring coverage with minimal set size. A more advanced technique, conditional CP (Cherian et al., 2025), dynamically adjusts coverage guarantees based on the difficulty of the input, optimizing prediction set size while maintaining reliability.
Consistency-Based Methods. Consistency-based uncertainty estimation methods analyze the agreement between multiple generated responses from an LLM to determine uncertainty. The underlying assumption is that if the model is confident, its responses should be consistent, while high variability among responses suggests uncertainty. One approach leverages Jaccard similarity (Lin et al., 2023), which measures the overlap between words in different generations. This method evaluates the deviation from self-consistency, where a high Jaccard similarity across generations implies low uncertainty.
However, in natural language generation (NLG) tasks, word-level similarity alone is insufficient, as different responses can convey the same meaning using different phrasing. To address this problem, some methods incorporate external models to assess semantic similarity rather than relying solely on lexical overlap.
3.4.3. Multiple Rounds Generation with External Models
Semantic-based uncertainty estimation methods extend multiple rounds of generation by incorporating external models to assess the consistency of generated responses beyond lexical similarity.
Semantic Entropy. Semantic Entropy (SE) (Kuhn et al., 2023), refines uncertainty estimation by clustering generated responses based on semantic equivalence. This approach uses a Natural Language Inference (NLI) model to determine entailment relationships among responses, grouping them into meaning-preserving clusters. Instead of calculating entropy over individual responses, SE computes entropy over these clusters. Another method, Kernel Language Entropy (KLE), applies a kernel-based framework to quantify semantic uncertainty (Nikitin et al., 2024). To enhance the performance of semantic entropy, KLE represents them in a semantic space using positive semidefinite kernels. By computing von Neumann entropy over these response distributions, KLE provides an even more fine-grained measure of uncertainty that considers nuanced semantic variations.
Semantic Similarity. Semantic similarity uncertainty methods use Natural Language Inference (NLI) models to measure the semantic relationships between multiple generated responses from an LLM (Lin et al., 2023). Instead of relying on lexical overlap, these approaches construct a similarity matrix based on entailment and contradiction scores between generated outputs. A confident model produces responses with high internal consistency, while greater semantic dispersion in the similarity matrix indicates higher uncertainty.
To quantify this dispersion, graph-based spectral metrics are applied. Eccentricity (Ecc) measures the spread of response variability, eigenvalue-based measures (Eig) analyze the spectral properties of the similarity matrix to detect uncertainty, and degree (Deg) evaluates response connectivity, with higher degrees indicating stronger confidence. Chen et al. (2025) proposes MD-UQ, which enhanced the performance of graph-based spectral metrics by using an additional knowledge dimension input extracted from an auxiliary LLM and tensor decomposition while Da et al. (2024a) also enhanced the performance by analyzing the directed graph instead of undirected graph originally from Lin et al. (2023).
4. Evaluation of Uncertainty in LLMs
4.1. Benchmark Datasets
Category Benchmarks
Reading Comprehension TriviaQA (Joshi et al., 2017), CoQA (Reddy et al., 2019), RACE (Lai et al., 2017), CosmosQA (Huang et al., 2019), SQuAD (Rajpurkar et al., 2016), HotpotQA (Yang et al., 2018)
Reasoning & Math StrategyQA (Geva et al., 2021), HotpotQA (Yang et al., 2018), GSM8K (Cobbe et al., 2021), CalibratedMath (Lin et al., 2022a)
Factuality TruthfulQA (Lin et al., 2022b), FEVER (Thorne et al., 2018), HaluEval (Li et al., 2023a), FActScore (Min et al., 2023)
General Knowledge MMLU (Hendrycks et al., 2020), GPQA (Rein et al., 2024), HellaSwag (Zellers et al., 2019)
Consistency & Ambiguity ParaRel (Elazar et al., 2021), AmbigQA (Min et al., 2020b), AmbigInst (Hou et al., 2024)
Table 3. Categorization of datasets and benchmarks.
Datasets used in previous studies can be organized into several categories based on their focus. Reading comprehension benchmarks include CoQA for conversational Q&A, RACE for general reading comprehension, TriviaQA for fact-based questions, CosmosQA for contextual understanding, SQuAD for question-answering on passages, and HotpotQA for multi-hop reasoning. Reasoning and math benchmarks include HotpotQA and StrategyQA, which test multi-hop reasoning, GSM8K for solving math problems, and CalibratedMath, designed to evaluate confidence expression in arithmetic. Factuality evaluation draws on datasets such as TruthfulQA for addressing common misconceptions, FEVER for claim verification, and HaluEval for detecting hallucinations, and annotated FActScore dataset for evaluating the factuality of long-form text generated by LLMs (Zhang et al., 2024a). Several general knowledge benchmarks have been adapted for uncertainty quantification, such as MMLU for a wide range of subjects, GPQA for multiple-choice questions in physical sciences, and HellaSwag for common-sense reasoning through sentence completion. These benchmarks can be adapted for UQ because the task reduces to a binary classification problem—determining whether the model is confident or uncertain. The structured nature of these benchmarks allows for clear evaluation of the model’s confidence in its predictions.
Additional categories include consistency benchmarks such as ParaRel, which tests semantic consistency across 328 paraphrases for 38 relations, and datasets like AmbigQA and AmbigInst, which feature inherent ambiguities. Ambiguity datasets are useful in uncertainty quantification (UQ) evaluation because they introduce aleatoric uncertainty by highlighting cases where multiple plausible interpretations exist, helping to assess how well models distinguish between data-driven randomness and model-based uncertainty. These datasets enable a more precise decomposition of uncertainty into aleatoric and epistemic components, improving model reliability and interpretability.
Recently, there have been efforts to introduce dedicated uncertainty quantification (UQ) benchmarks for large language models (LLMs). Yang et al. (2024) introduced MAQA, a dataset specifically designed to evaluate epistemic uncertainty in language models. Fadeeva et al. (2023) created LM-Polygraph, which was later adopted as a comprehensive uncertainty benchmark by Vashurin et al. (2024). Ye et al. (2025) developed a benchmark using conformal prediction across five common NLP tasks. These contributions represent specialized datasets explicitly designed to assess UQ capabilities in LLMs, rather than adapting existing general-purpose benchmarks. Overall, we show the categorization of datasets and benchmarks for UQ in Table 3.
4.2. Evaluation Metrics
Uncertainty quantification (UQ) is often evaluated as a binary classification task, the rationale being that high uncertainty should correspond to low expected accuracy. This is typically modeled by assigning a binary label to each response with a correctness function and using the uncertainty estimates to predict the label. AUROC (Area Under the Receiver Operating Characteristic curve), which measures how effectively the uncertainty score separates correct from incorrect responses, is often used. With values ranging from 0 to 1, higher AUROCs indicate better performance. Responses with confidence above the threshold are classified as predicted positives, while those below are treated as predicted negatives. Many prior studies use AUROC to evaluate how well uncertainty score discriminates correct from incorrect predictions (Chen and Mueller, 2024; Lin et al., 2023; Kuhn et al., 2023; Xiong et al., 2024a; Liu et al., 2024c).
Similarly, AUPRC (Area Under the Precision-Recall Curve) and AUARC (Area Under the Accuracy-Rejection Curve) (Nadeem et al., 2009) also offer further insights into uncertainty quantification. AUPRC measures how well the uncertainty score separates correct from incorrect responses (Ling et al., 2024), while AUARC assesses how effectively the uncertainty measure aids in selecting accurate responses by determining which uncertain questions to reject (Lin et al., 2024).
In the context of NLG where the correctness label is hard to obtain, researchers also compute heuristic-based (fuzzy matching) metrics such as BLEU (Papineni et al., 2002) and ROUGE (Kuhn et al., 2023) between the generated text and the reference output(s) to gauge the quality. However, these metrics often fail to capture semantic fidelity or factual correctness. Consequently, many researchers are increasingly turning to LLM-as-a-judge evaluations, wherein a large language model (e.g., GPT-4) is prompted to assess text quality or correctness. This approach can capture nuanced aspects like coherence, style, and factuality, but also introduces risks of bias and inconsistency. Human annotation, however, is expensive and is often limited to a small scale (Zhang et al., 2024a; Kuhn et al., 2023).
Apart from the binary classification framework, there are also multiple evaluation methods designed for specific treatment of uncertainty, sometimes qualitative. For example, focusing on decomposing aleatoric and epistemic uncertainty, Hou et al. (2024) evaluates only the aleatoric part by using AmbigQA (Min et al., 2020b), as high ambiguity questions should incur higher aleatoric uncertainty (whereas math questions, for examples, might have lower). The evaluation in Giulianelli et al. (2023), on the other hand, is a comparison between the variability of human production (generation) with that of the LM. With an emphasis on UQ for longer generations, Zhang et al. (2024a) compare the uncertainty estimate against FActScore (Min et al., 2023), as the “correctness” of a long paragraph could be ill-defined or ambiguous.
5. Applications in LLMs
LLMs are increasingly applied in diverse domains, offering flexibility and reasoning capabilities (Da et al., 2025b). However, uncertainty quantification (UQ) is crucial for ensuring their reliability, particularly in high-stakes applications like Education (Lee et al., 2024) and healthcare (Savage et al., 2025; Da et al., 2024d). This section will introduce the applications that integrate the UQ of LLMs from different domains.
Robotics. LLM-based robotic planning suffers from ambiguity and hallucinations, motivating the need for uncertainty quantification in the planning loop. For example, closed-loop planners (Zheng et al., 2024) employs an uncertainty-based failure detector to continuously assess and adjust plans in real-time, while non-parametric UQ method (Tsai et al., 2024) use an efficient querying strategy to improve reliability. In addition, ’LAP’ (Mullen Jr and Manocha, 2024) integrates action feasibility checks to align the LLM’s confidence with real-world constraints, improving success rates from approximately 70% to 80%. Similarly, adaptive skill selection (Ong et al., 2024) dynamically adjusts thresholds for alternative paths, achieving higher success rates in the ALFRED domain. Finally, introspective planning (Liang et al., 2024b) enables LLMs to self-assess uncertainty, enhancing safety and human-robot collaboration.
Transportation. Preliminary research explores how LLMs can enhance transportation systems (Da et al., 2024c). For example, LLM inference has been used to bridge the sim-to-real gap in Traffic Signal Control policies (Da et al., 2023b, a, 2024b), and LLM agents have been shown to help smooth mixed-autonomy traffic (Yao et al., 2024). However, both cases reveal the potential risk posed by hallucination. A few works have investigated the uncertainty measure while using the LLMs as in (de Zarzà et al., 2023); it tries to link the use of visual language models (VLMs) with deep probabilistic programming for uncertainty quantification while conducting multimodal traffic accident forecasting tasks and (Da et al., 2024a) shows a general solution by modeling the responses as a directed-graph and captures the response uncertainty. Further, UQ exploration anchored on traffic-related LLM use is expected.
Healthcare. In healthcare, LLMs (VLMs) can be good reference for diagnosis, but uncertainty is a critical dimension that should be output together with the generation for more reliable treatment plans. Such as (Savage et al., 2024) evaluates the ability of uncertainty metrics to quantify LLM confidence when performing diagnosis and treatment selection tasks by assessing the properties of discrimination and calibration.  (Chen et al., 2024) first quantify uncertainty in white-box models with access to model parameters and output logits, and by this, it reveals that an effective reduction of model uncertainty can be achieved by using the proposed multi-tasking and ensemble methods in EHRs. With these findings, however, as  (Wu et al., 2024) benchmarks popular uncertain quantification methods with different model sizes on medical question-answering datasets, authors find that the challenge of UQ for medical applications is still severe.
Education. Many LLM-related products are popular to guide the thinking of students in AI for Education research (Zhang et al., 2024e). However, it is critical to ensure the trustworthiness of LLMs if to participate in a person’s thinking logic learning process (Razafinirina et al., 2024). The work in (Tziakouri, 2024) discusses the solutions of black-box models UQ evaluation, including Graph Laplacian Eigenvalue Sum, Lexical Similarity, and their reliability. (Da et al., 2025a) proposes a method that elicits the reasoning topology of the LLM answering process, which can guide the step-by-step analysis of the LLM’s reasoning steps before applying it to the student reference guide.
Conclusion. There are many different fields that discuss the need for uncertainty quantification during the LLMs execution process, and the above-discussed ones are only major aspects for now, the multi-agent energy management, operation research, etc, all employ LLMs and would require such consideration respectively.
6. Challenges and Future Directions
While significant strides have been made in integrating uncertainty quantification into Large Language Models, several unaddressed challenges persist. This section will explore these unresolved issues, ranging from efficiency-performance trade-offs to cross-modal uncertainty, and outline promising avenues for future research, aiming to advance the reliability and trustworthiness of LLMs in high-stakes applications.
Efficiency-Performance Trade-offs. Multi-sample uncertainty methods incur prohibitive costs for trillion-parameter LLMs (\$12k per million queries (Li et al., 2024a)), yet yield marginal reliability gains ($\leq 0.02$ AUROC improvement (Xiong et al., 2024b)). Hybrid approaches combining low-cost proxies (attention variance (Heo et al., 2018), hidden state clustering (Nikitin et al., 2024)) could resolve this by achieving 90% of maximal performance at 10% computational cost. For example, precomputing uncertainty “hotspots” during inference could trigger targeted multi-sampling only for high-risk outputs like medical diagnoses.
Interpretability Deficits. Users struggle to distinguish whether uncertainty stems from ambiguous inputs (“Explain quantum gravity”), knowledge gaps, or decoding stochasticity—a critical barrier in high-stakes domains. Modular architectures that decouple uncertainty estimation layers (Huang et al., 2025b; Sensoy et al., 2018) or employ causal tracing of transformer attention pathways (Wang et al., 2024b) could clarify uncertainty origins. For instance, perturbing model weights (Gal and Ghahramani, 2016) might reveal parametric uncertainty in low-resource languages, while input modules flag underspecified terms (”this country”) for clarification.
Cross-Modality Uncertainty. Integrating vision, text, and sensor data introduces misaligned confidence estimates between modalities: LVLMs exhibit 2.4×2.4$\times$2.4 × higher uncertainty in visual vs. textual components (Zhang et al., 2024b), causing 63% of errors in multi-modal QA (Zhang et al., 2024d). Dynamic contrastive decoding and uncertainty-aware fusion protocols show promise(Huo et al., 2024; Suo et al., 2025), but require domain-specific adaptations (e.g., aligning radiology images with reports (Li et al., 2023b; Zhao et al., 2023b)). Future work must develop unified uncertainty embeddings to harmonize modality confidence scales and adversarial training against cross-modal backdoor attacks (Zhang et al., 2024c; Walmer et al., 2022; Liang et al., 2024a).
Interventions for Uncertainty. Real-time uncertainty monitoring enables self-correction (e.g., retracting erroneous reasoning steps (Yin et al., 2024)), but adds 300 to 500ms latency per decoding iteration (Xiong et al., 2024b). Lightweight monitors via attention variance distillation (Gong et al., 2021; Huang et al., 2022) and human-in-the-loop calibration (Wang et al., 2022) could mitigate overhead while addressing security risks like adversarial uncertainty manipulation (100% attack success rates (Zeng et al., 2024)). Certified defense mechanisms and causal propagation tracing (Goyal et al., 2023) are critical for deploying uncertainty-guided interventions in sensitive applications.
UQ Evaluation. There are several challenges to effectively evaluating UQ quality. There are limitations in adopting a binary classification evaluation approach: It is challenging to decide if a free-form generation is correct even in a QA dataset, let alone more open-ended tasks. Also, even LLM-as-a-judge suffers systematic biases including self-preference bias (where an LLM judge may favor answers written in a style similar to its own) (Panickssery et al., 2024), length bias (where LLM judges often prefer longer, more detailed answers) (Lin et al., 2022b), and position bias (where the order of presented answers influences the judgment) (Zheng et al., 2023b). Moreover, current evaluation methods, such as AUROC and AUARC, fall short in capturing nuanced, “meaningful” uncertainty. For instance, they often do not differentiate between cases where a model is confidently wrong (hallucinations) and cases where it expresses a prudent level of uncertainty. Finally, a significant practical challenge in evaluating UQ for LLMs is the lack of suitable datasets that align with what we want to measure. Many benchmarks exist to test LLM capabilities (from reading comprehension to factual QA), but few are explicitly designed for UQ.
7. Conclusion
In this survey, we offer a comprehensive overview of uncertainty quantification (UQ) in Large Language Models (LLMs). Initially, we introduce the fundamental concepts relevant to both UQ and LLMs, highlighting the importance of reliability in high-stakes applications. Following this, we propose a detailed taxonomy for characterizing uncertainty dimensions in LLMs, including input, reasoning, parameter, and prediction uncertainty. In terms of methodologies, we systematically evaluate UQ methods using our novel taxonomy, thoroughly reviewing their effectiveness across different uncertainty types. Ultimately, we identify and discuss some of the persistent challenges in UQ for LLMs, providing insightful directions for future research. The primary goal of this survey is to promote a more seamless integration of UQ techniques into LLM development, motivating both machine learning researchers and practitioners to delve into this rapidly advancing area.